# Polarization of reflected X-ray emission from Sgr A molecular complex: multiple flares, multiple sources?

### v1

@@ 6 Discussion
While confirming that the total polarization signal from the large region centred on the Sgr A complex might come from Sgr A*, we found a significantly different polarization signal for the brightest and highly time-variable central part of it.
@@
We consider two options that, within uncertainties, are consistent with the data. Option 1: The signal from the compact bright region has the same polarization direction as in the large region, but a significantly lower PD. In this case, accumulated statistics still allow significant fluctuations of the PA, even though the probability of this is low. Option 2: The PA is genuinely different from that inferred for the large region.
@@
The former case leaves open a possibility of Sgr A* being the primary source of illumination, but implies that the brightest reflection spot is located at another distance (lower PD implies larger distance along the line of sight), lending further support to the hypothesis of several flaring episodes (as already suggested in Clavel et al., 2013, based on the spatial and temporal variability patterns).
@@
The latter case can be reconciled with a Sgr A* flare only if its primary emission was strongly polarized with the sky projection of the electric field vector aligned with the source-cloud direction (as shown in the bottom panel of Figure 2 and in Figure 4 of Khabibullin et al., 2020a). This configuration is possible only when the cloud is located at the far side of the reflecting paraboloid, allowing it to ‘mirror’ the intrinsic polarization of the primary source. Although infrared emission of the regular Sgr A* flares is known to be polarized (Eckart et al., 2006; Shahzamanian et al., 2015), and the spatially resolved emission from the currently present accretion flow has been measured (e.g., Event Horizon Telescope Collaboration et al., 2024), the possibility of extrapolating these measurements to a hypothetical orders-of-magnitude brighter accretion episode (e.g., Dovčiak et al., 2008), which might be responsible for the observed X-ray echo, remains uncertain.
@@
At the same time, the spectral analysis of the variable diffuse emission from both large and the brightest regions indicates that the spectral shapes of the reflected emission are rather similar, with the slight variations in the equivalent width of the fluorescent line of neutral iron well within the possible effects of different scattering angle or metallicity of the reflecting cloud. This appears to disfavor intrinsic polarization of the primary emission since, in this case, apparent rotation of the polarization vector would be associated with substantial variations in the ratio of the fluorescent line to the reflected continuum. In other words, the similarity of the observed spectral shape indicates that the scattering angle in both cases should be similar as well. Taken at face value, this would indicate the presence of at least two primary sources responsible for the current X-ray echo(es) propagating through the Sgr A complex. While for the large region Sgr A* remains a very viable candidate, the polarization vector of the brightest region is perpendicular to the direction towards the Arches stellar cluster, which is one of the youngest massive clusters in the Milky Way, and is known to host highly variable compact sources.
@@
Thus, we come to a picture of a plausible ‘global’ illumination of the large molecular complexes in the CMZ by a luminous flare in its centre, and the presence of ‘local’ time–and space-variable reflection signal(s). Although such a picture means a great increase in complexity compared to the simplest ‘single source, single (short) flare’ scenario, it appears to be a possible alternative.
@@
Contrary to the case of Sgr A* being a primary illumination source, it is more difficult to estimate the required fluence of the initial X-ray flare because of the (more) uncertain relative distance between the primary source and the reflector. In particular, for a source located close to the reflecting cloud, the required fluence of the flare will be even smaller than the value inferred for Sgr B2 or Sgr A complexes in the Sgr A* scenario. This opens the possibility for even less dramatic transient events to be responsible for the observed ‘local’ reflection echoes. Further monitoring of the morphological and spectral changes (e.g., Ponti et al., 2010; Churazov et al., 2017c), in combination with the data on the molecular gas distribution in velocity space (e.g., Ponti et al., 2013; Clavel et al., 2013; Churazov et al., 2017b; Khabibullin et al., 2020a; Brunker et al., 2025), should help us to pinpoint the location of the primary source with more confidence, and, as a result, better quantify required energetics of the flare and its possible nature.
@@ 7 Conclusions
We report the results of two deep observations of the currently brightest (in reflected emission) molecular complex Sgr A taken with the imaging X-ray polarimeter IXPE in 2022 and 2023, and supplement them with the data of a quasi-simultaneous Chandra monitoring campaign of the same region. We confirm the previous polarization measurement for a large region encompassing the Sgr A complex with high () significance, but also reveal an inconsistent polarization pattern for the brightest reflection region in its center.
@@
This compact region has a smaller PD compared to the large region, and a PA that is almost perpendicular to the expected direction in the Sgr A* unpolarized-flare scenario. While the statistical significance of this result is at the level of less than 3, it might indicate that several illumination fronts have propagated through the CMZ simultaneously, with at least one of them not due to a Sgr A* flare. The scenario of Sgr A* being the primary source of illumination for both regions is still possible, but requires that either (i) currently the brightest compact region of the Sgr A complex is located at a significantly farther distance (and hence is illuminated by an older flare), or (ii) that the primary emission was itself strongly polarized in a manner allowing almost perpendicular rotation of the polarization vector relative to the unpolarized case (e.g., in the configuration illustrated in Khabibullin et al., 2020a). In fact, the latter option also requires the reflecting medium being well ‘behind’ Sgr A*.
@@
If instead the inferred polarization direction is indeed determined by direction to the primary source, it could be associated with the Arches stellar cluster or a currently unknown transient source located in a closer vicinity to the illuminated cloud (see a discussion of implications of such a scenario in Sunyaev & Churazov, 1998). Although a scenario of multiple (transient) illuminating sources is straightforward to model (e.g., Molaro et al., 2014), the arising complexity of spectral variability and polarization patterns requires sensitive monitoring observations of many individual reflecting complexes (e.g., Marin et al., 2014).
@@
Significantly deeper observations with IXPE would be required to unequivocally distinguish between the scenarios. Given the consistency of the observed noise levels with expected statistical fluctuations, accumulation of an additional 1.1 Ms of clean exposure time would be needed to ensure the confidence level detection for the brightest reflection region (if the real signal is indeed at the detection level and stays constant in time). A combination with continued high-resolution imaging, micro-calorimetric spectroscopy (e.g., with XRISM observatory, Tashiro et al., 2025), and global monitoring of the new reflection regions via CMZ-wide surveys (e.g., Khabibullin et al., 2020a; Semena et al., 2024) offer a promising path forward (along with the roadmap outlined in Churazov et al., 2019). Next-generation X-ray polarimetry missions will be required to disentangle the full complexity of X-ray reflection signals if the scenario of multiple sources and/or multiple flares is confirmed. An additional window might be offered by measuring polarization and spectra at higher energies, ¿10 keV, where the reflected continuum should be dominant over other emission components. Prospective missions such as the Enhanced X-ray Polarimetry Observatory (EXPO), evolution of the NGXP concept (Soffitta et al., 2021), would be ideally suited for this task.
Acknowledgements
@@
The Imaging X-ray Polarimetry Explorer (IXPE) is a joint US and Italian mission. The US contribution is supported by the National Aeronautics and Space Administration (NASA) and led and managed by its Marshall Space Flight Center (MSFC), with industry partner Ball Aerospace (contract NNM15AA18C). The Italian contribution is supported by the Italian Space Agency (Agenzia Spaziale Italiana, ASI) through contract ASI-OHBI-2022-13-I.0, agreements ASI-INAF-2022-19-HH.0 and ASI-INFN-2017.13-H0, and its Space Science Data Center (SSDC) with agreements ASI-INAF-2022-14-HH.0 and ASI-INFN 2021-43-HH.0, and by the Istituto Nazionale di Astrofisica (INAF) and the Istituto Nazionale di Fisica Nucleare (INFN) in Italy. This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC).
We thank Dylan Maurel of the IXPE Science Operations Center for manually reprocessing the IXPE data.
IK acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. JS thanks GACR project 21-06825X. AV, RK, and WF acknowledge support from the Smithsonian Institution and the Chandra High Resolution Camera Project through NASA contract NAS8-03060. IL was funded by the European Union ERC-2022-STG – BOOTES – 101076343. RF, ECo, ADM, PSo, FLM, FMu and SF are partially supported by MAECI with grant CN24GR08 “GRBAXP: Guangxi-Rome Bilateral Agreement for X-ray Polarimetry in Astrophysics”.
Appendix A Systematic uncertainties

### v2

@@ 6 Discussion
While confirming that the total polarization signal from the large region centred on the Sgr A complex might come from Sgr A*, we found a significantly different polarization signal for the brightest and highly time-variable central part of it. We considered two options which, within uncertainties, are consistent with the data. In one option, the signal from the compact bright region has the same polarization direction as in the large region but a significantly lower PD. In this case, accumulated statistics still allow for significant fluctuations of the PA, even though the probability of this is low. In the second option, the PA is genuinely different from that inferred for the large region.
@@
The first case leaves open a possibility of Sgr A* being the primary source of illumination, but it implies that the brightest reflection spot is located at another distance (a lower PD implies larger distance along the line of sight), lending further support to the hypothesis of several flaring episodes (as already suggested in Clavel et al., 2013; Chuard et al., 2018, based on the spatial and temporal variability patterns in different molecular complexes).
@@
The latter case can be reconciled with a Sgr A* flare only if its primary emission is strongly polarized with the sky projection of the electric field vector aligned with the source-cloud direction (as shown in the bottom panel of Fig. 2 and in Fig. 4 of Khabibullin et al., 2020b). This configuration is possible only when the cloud is located at the far side of the reflecting paraboloid, which allows it to ”mirror” the intrinsic polarization of the primary source. Although infrared emission of the regular Sgr A* flares is known to be polarized (Eckart et al., 2006; Shahzamanian et al., 2015; Gravity Collaboration et al., 2023) and the spatially resolved emission from the currently present accretion flow has been measured (e.g., Event Horizon Telescope Collaboration et al., 2024), the possibility of extrapolating these measurements to a hypothetical orders-of-magnitude brighter accretion episode (e.g., Dovčiak et al., 2008), which might be responsible for the observed X-ray echo, remains uncertain.
@@
At the same time, the spectral analysis of the variable diffuse emission from both large and the brightest regions indicates that the spectral shapes of the reflected emission are rather similar, with the slight variations in the equivalent width of the fluorescent line of neutral iron well within the possible effects of different scattering angle or metallicity of the reflecting cloud. This appears to disfavor intrinsic polarization of the primary emission since, in this case, apparent rotation of the polarization vector would be associated with substantial variations in the ratio of the fluorescent line to the reflected continuum. In other words, the similarity of the observed spectral shape indicates that the scattering angle in both cases should be similar as well.
Taken at face value, this would indicate the presence of at least two primary sources responsible for the current X-ray echo(es) propagating through the Sgr A complex. While for the large region Sgr A* remains a very viable candidate, the polarization vector of the brightest region is perpendicular to the direction toward the Arches stellar cluster, which is one of the youngest massive clusters in the Milky Way, and is known to host highly variable compact sources.
@@
Thus, we come to a picture of a plausible ”global” illumination of the large molecular complexes in the CMZ by a luminous flare in its centre, and the presence of ”local” time–and space-variable reflection signal(s). Although such a picture means a great increase in complexity compared to the simplest ”single source, single (short) flare” scenario, it appears to be a possible alternative.
@@
Contrary to the case of Sgr A* being a primary illumination source, it is more difficult to estimate the required fluence of the initial X-ray flare because of the (more) uncertain relative distance between the primary source and the reflector. In particular, for a source located close to the reflecting cloud, the required fluence of the flare will be smaller than the value inferred for Sgr B2 or Sgr A complexes in the Sgr A* scenario. This opens the possibility for even less dramatic transient events to be responsible for the observed ‘local’ reflection echoes.
For instance, for a cloud located at , where is the projected distance from Sgr A* in the sky plane, i.e., farther away from us than Sgr A*, the degree of polarization will be (see Eq. (6)). For the brightest patch, . Similarly to the case discussed above, this scenario requires the scattering cloud located farther away from us than Sgr A*. In this case, the scattering angle can be close to 180 degrees, and the scattering by itself does not induce any strong polarization. Instead, the reflected emission will have the degree and orientation of polarization corresponding to the incident emission from the primary source. Among these potential sources of the illuminating X-ray flux, the Quintuplet cluster has the smallest projected distance to the brightest patch. This lowers the minimum required luminosity by an order of magnitude compared to the Sgr A* case (assuming the Quintuplet and the bright patch are at the same distance from us). However, the shorter light crossing time implies that the primary source in the Quintuplet was active some 25 years ago. Adjusting the geometrical parameters, one can shift the flare to years ago without a strong increase in the required luminosity. This scenario cannot be unambiguously excluded.
Further monitoring of the morphological and spectral changes (e.g., Ponti et al., 2010; Churazov et al., 2017c), in combination with the data on the molecular gas distribution in velocity space (e.g., Ponti et al., 2013; Clavel et al., 2013; Churazov et al., 2017b; Khabibullin et al., 2020b; Brunker et al., 2025), should help us to pinpoint the location of the primary source with more confidence, and, as a result, better quantify required energetics of the flare and its possible nature.
@@ 7 Conclusions
We have reported the results of two deep observations of the currently brightest (in reflected emission) molecular complex Sgr A taken with the imaging X-ray polarimeter IXPE in 2022 and 2023, and we supplemented them with the data of a quasi-simultaneous Chandra monitoring campaign of the same region. We confirm the previous polarization measurement for a large region encompassing the Sgr A complex with high () significance, but we also revealed an inconsistent polarization pattern for the brightest reflection region in its center.
@@
This compact region has a smaller PD compared to the large region, and the PA is almost perpendicular to the expected direction in the Sgr A* unpolarized-flare scenario. While the statistical significance of this result is at the level of less than 3, it might indicate that several illumination fronts have propagated through the CMZ simultaneously, with at least one of them not being due to a Sgr A* flare. The scenario of Sgr A* as the primary source of illumination for both regions is still possible, but it requires either (i) that the currently brightest compact region of the Sgr A complex is located at a significantly farther distance (and hence is illuminated by an older flare) or (ii) that the primary emission was itself strongly polarized in a manner allowing almost perpendicular rotation of the polarization vector relative to the unpolarized case (e.g., in the configuration illustrated in Khabibullin et al., 2020b). In fact, the latter option also requires the reflecting medium to be well ”behind” Sgr A*.
@@
If instead the inferred polarization direction is indeed determined by the direction to the primary source, it could be associated with the Arches stellar cluster or a currently unknown transient source located in a closer vicinity to the illuminated cloud (see a discussion of implications of such a scenario in Sunyaev and Churazov, 1998). Although a scenario of multiple (transient) illuminating sources is straightforward to model (e.g., Molaro et al., 2014), the arising complexity of spectral variability and polarization patterns requires sensitive monitoring observations of many individual reflecting complexes (e.g., Marin et al., 2014).
Significantly deeper observations with IXPE would be required to unequivocally distinguish between the scenarios. Given the consistency of the observed noise levels with the expected statistical fluctuations, accumulation of an additional 1.1 Ms of clean exposure time would be needed to ensure the confidence level detection for the brightest reflection region (if the real signal is indeed at the detection level and would stay constant in time). There is, however, an indication of a decrease in the observed surface brightness from the Sgr A complex, which makes this task even more challenging.
@@
A combination of continued high-resolution imaging, micro-calorimetric spectroscopy (e.g., with XRISM observatory, Tashiro et al., 2025), and global monitoring of the new reflection regions via CMZ-wide surveys (e.g., Khabibullin et al., 2020b; Semena et al., 2024) offers a promising path forward (along with the roadmap outlined in Churazov et al., 2019). Next-generation X-ray polarimetry missions will be required to disentangle the full complexity of X-ray reflection signals if the scenario of multiple sources and/or multiple flares is confirmed. An additional window might be offered by measuring polarization and spectra at higher energies (¿10 keV), where the reflected continuum should be dominant over other emission components. Prospective missions such as the Enhanced X-ray Polarimetry Observatory, an evolution of the NGXP concept (Soffitta et al., 2021), would be ideally suited for this task.
Acknowledgements.
@@
The Imaging X-ray Polarimetry Explorer (IXPE) is a joint US and Italian mission. The US contribution is supported by the National Aeronautics and Space Administration (NASA) and led and managed by its Marshall Space Flight Center (MSFC), with industry partner Ball Aerospace (contract NNM15AA18C). The Italian contribution is supported by the Italian Space Agency (Agenzia Spaziale Italiana, ASI) through contract ASI-OHBI-2022-13-I.0, agreements ASI-INAF-2022-19-HH.0 and ASI-INFN-2017.13-H0, and its Space Science Data Center (SSDC) with agreements ASI-INAF-2022-14-HH.0 and ASI-INFN 2021-43-HH.0, and by the Istituto Nazionale di Astrofisica (INAF) and the Istituto Nazionale di Fisica Nucleare (INFN) in Italy. This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC). We thank Dylan Maurel of the IXPE Science Operations Center for manually reprocessing the IXPE data.IK acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. JS thanks GACR project 21-06825X. AV, RK, and WF acknowledge support from the Smithsonian Institution and the Chandra High Resolution Camera Project through NASA contract NAS8-03060. IL was funded by the European Union ERC-2022-STG – BOOTES – 101076343. RF, ECo, ADM, PSo, FLM, FMu, and SF are partially supported by MAECI with grant CN24GR08 “GRBAXP: Guangxi-Rome Bilateral Agreement for X-ray Polarimetry in Astrophysics”. We are grateful to the anonymous referee for a constructive report that helped us to improve the paper.References
- 3D MC. II. X-Ray Echoes Reveal a Clumpy Molecular Cloud in the CMZ. AJ 169 (4), pp. 213. External Links: Document, 2501.07669, ADS entry Cited by: §1, §4.
- Disentangling the Galactic centre X-ray reflection signal using XMM-Newton data. A&A 698, pp. A313. External Links: Document, 2505.09672, ADS entry Cited by: §4.3, §6.
- XSPEC: The First Ten Years. In Astronomical Data Analysis Software and Systems V, G. H. Jacoby and J. Barnes (Eds.), ASP Conf. Ser. January 17 ADS entry Provided by the SAO/NASA Astrophysics Data System @inproceedings{1996ASPC..101…17A, author = {{Arnaud}, K.~A.}, title = {{XSPEC: The First Ten Years}}, booktitle = {Astronomical Data Analysis Software and Systems V}, year = {1996}, editor = {{Jacoby}, George H. and {Barnes}, Jeannette}, series = {ASP Conf. Ser. %Astronomical Society of the Pacific Conference Series}, volume = {101}, month = {January}, pages = {17}, adsurl = {https://ui.adsabs.harvard.edu/abs/1996ASPC..101…17A}, adsnote = {Provided by the SAO/NASA Astrophysics Data System}} , Vol. 101. Cited by: §4.3.
- 3D MC. I. X-Ray Tomography Begins to Unravel the 3D Structure of a Molecular Cloud in our Galaxy’s Center. ApJL 982 (1), pp. L20. External Links: Document, 2501.07717, ADS entry Cited by: §1, §1, §6.
- Discovery of X-ray flaring activity in the Arches cluster. A&A 525, pp. L2. External Links: Document, 1011.4062, ADS entry Cited by: §4.2.
- The X-ray lightcurve of Sagittarius A* over the past 150 years inferred from Fe-K line reverberation in Galactic centre molecular clouds. A&A 545, pp. A35. External Links: Document, 1207.1436, ADS entry Cited by: §1.
- Glimpses of the past activity of Sgr A★ inferred from X-ray echoes in Sgr C. A&A 610, pp. A34. External Links: Document, 1712.02678, ADS entry Cited by: §1, §6, §6.
- Polarization and long-term variability of Sgr A* X-ray echo. MNRAS 468 (1), pp. 165–179. External Links: Document, 1612.00180, ADS entry Cited by: §1, §1, §4.3, §4, §6, footnote 9.
- Can Sgr A* flares reveal the molecular gas density PDF?. MNRAS 471 (3), pp. 3293–3304. External Links: Document, 1705.05878, ADS entry Cited by: §1, §2, §6.
- Not that long time ago in the nearest galaxy: 3D slice of molecular gas revealed by a 110 yr old flare of Sgr A*. MNRAS 465 (1), pp. 45–53. External Links: Document, 1610.06699, ADS entry Cited by: §1, §4, §4, §6.
- Polarization of X-ray emission from the Sgr B2 cloud. MNRAS 330 (4), pp. 817–820. External Links: Document, astro-ph/0111065, ADS entry Cited by: §1, §1, §2, §4.3, footnote 8.
- Pulsar-wind-nebula-powered Galactic center X-ray filament G0.13-0.11. Proof of the synchrotron nature by IXPE. A&A 686, pp. A14. External Links: Document, 2312.04421, ADS entry Cited by: §1, §5.2, §5.2.
- Probing 3D Density and Velocity Fields of ISM in Centers of Galaxies with Future X-Ray Observations. BAAS 51 (3), pp. 325. External Links: Document, 1903.06429, ADS entry Cited by: §7.
- Variation of the X-ray non-thermal emission in the Arches cloud.. MNRAS 443, pp. L129–L133. External Links: Document, 1406.5727, ADS entry Cited by: §4.2.
- Echoes of multiple outbursts of Sagittarius A⋆ revealed by Chandra. A&A 558, pp. A32. External Links: Document, 1307.3954, ADS entry Cited by: §1, §4, §6, §6.
- Les auréoles lumineuses des Novæ. AnAp 2, pp. 271. External Links: ADS entry Cited by: §1.
- Prospects for IXPE and eXTP polarimetric archaeology of the reflection nebulae in the Galactic center. A&A 643, pp. A52. External Links: Document, 2008.07830, ADS entry Cited by: §1.
- Handling the Background in IXPE Polarimetric Data. AJ 165 (4), pp. 143. External Links: Document, 2302.02927, ADS entry Cited by: §3.1, §5.2.
- Evidence for X-Ray Synchrotron Emission from Simultaneous Mid-Infrared to X-Ray Observations of a Strong Sgr A* Flare. ApJ 698 (1), pp. 676–692. External Links: Document, 0903.3416, ADS entry Cited by: §1.
- Thermal disc emission from a rotating black hole: X-ray polarization signatures. MNRAS 391 (1), pp. 32–38. External Links: Document, 0809.0418, ADS entry Cited by: §6.
- First simultaneous NIR/X-ray detection of a flare from Sgr A*. A&A 427, pp. 1–11. External Links: Document, astro-ph/0403577, ADS entry Cited by: §1.
- Polarimetry of near-infrared flares from Sagittarius A*. A&A 455 (1), pp. 1–10. External Links: Document, astro-ph/0610103, ADS entry Cited by: §6.
- First Sagittarius A* Event Horizon Telescope Results. VII. Polarization of the Ring. ApJ 964 (2), pp. L25. External Links: Document, ADS entry Cited by: §6.
- Prospects for a polarimetric mapping of the Sgr A molecular cloud complex with IXPE. A&A 655, pp. A108. External Links: Document, 2109.06678, ADS entry Cited by: §1.
- Powerful Radio Sources in Clusters of Galaxies - the Origin of Jet like Structures. SvAL 13, pp. 233. External Links: ADS entry Cited by: §2.
- Polarimetry and astrometry of NIR flares as event horizon scale, dynamical probes for the mass of Sgr A*. A&A 677, pp. L10. External Links: Document, 2307.11821, ADS entry Cited by: §6.
- Chandra Spectral and Timing Analysis of Sgr A*’s Brightest X-Ray Flares. ApJ 886 (2), pp. 96. External Links: Document, 1908.01781, ADS entry Cited by: §1.
- Time Variability of the Neutral Iron Lines from the Sagittarius B2 Region and Its Implication of a Past Outburst of Sagittarius A. PASJ 61, pp. S241. External Links: Document, ADS entry Cited by: §1.
- Über die Bewegung der Nebel in der Umgebung von Nova Persei. AN 157 (12), pp. 201. External Links: Document, ADS entry Cited by: §1.
- X-raying molecular clouds with a short flare: probing statistics of gas density and velocity fields. MNRAS 495 (1), pp. 1414–1432. External Links: Document, 1906.11579, ADS entry Cited by: §1, §2.
- Impact of intrinsic polarization of Sgr A* historical flares on (polarization) properties of their X-ray echoes. MNRAS 498 (3), pp. 4379–4385. External Links: Document, 2004.01960, ADS entry Cited by: §1, §4.3, §6, §6, §7, §7.
- SRG/eROSITA view of X-ray reflection in the Central Molecular Zone: a snapshot in September-October 2019. MNRAS 509 (4), pp. 6068–6076. External Links: Document, 2111.07899, ADS entry Cited by: §1, §1, §6, footnote 8.
- Analyzing the data from X-ray polarimeters with Stokes parameters. APh 68, pp. 45–51. External Links: Document, 1409.6214, ADS entry Cited by: §3.1.
- ASCA View of Our Galactic Center: Remains of Past Activities in X-Rays?. PASJ 48, pp. 249–255. External Links: Document, ADS entry Cited by: §1.
- First Hard X-Ray Detection of the Non-thermal Emission around the Arches Cluster: Morphology and Spectral Studies with NuSTAR. ApJ 781 (2), pp. 107. External Links: Document, 1312.2635, ADS entry Cited by: §4.2.
- NuSTAR and XMM-Newton observations of the Arches cluster in 2015: fading hard X-ray emission from the molecular cloud. MNRAS 468 (3), pp. 2822–2835. External Links: Document, 1612.03320, ADS entry Cited by: §1, §4.2.
- Investigating the origin of the faint non-thermal emission of the Arches cluster using the 2015-2016 NuSTAR and XMM-Newton X-ray observations. MNRAS 484 (2), pp. 1627–1636. External Links: Document, 1901.03121, ADS entry Cited by: §4.2.
- Sgr B2 hard X-ray emission with INTEGRAL after 2009: still detectable?. MNRAS 509 (2), pp. 1605–1613. External Links: Document, 2110.07401, ADS entry Cited by: §1, footnote 8.
- Prospects of 3D mapping of the Galactic Centre clouds with X-ray polarimetry. MNRAS 441 (4), pp. 3170–3176. External Links: Document, 1405.0898, ADS entry Cited by: §1, §1, §7.
- Reflection nebulae in the Galactic center: soft X-ray imaging polarimetry. A&A 576, pp. A19. External Links: Document, 1502.04894, ADS entry Cited by: §1, §1.
- X-ray polarization evidence for a 200-year-old flare of Sgr A∗. Natur 619 (7968), pp. 41–45. External Links: Document, 2304.06967, ADS entry Cited by: §1, §2, §3.1, §3.1, §4.3, §4.3, §5.2, §5.2, §5.3, footnote 9.
- Two sources of diffuse X-ray emission from the Galactic Centre. Nature 364 (6432), pp. 40–42. External Links: Document, ADS entry Cited by: §1.
- A thin diffuse component of the Galactic ridge X-ray emission and heating of the interstellar medium contributed by the radiation of Galactic X-ray binaries. A&A 564, pp. A107. External Links: Document, 1312.7029, ADS entry Cited by: §7.
- Probing the clumping structure of giant molecular clouds through the spectrum, polarisation and morphology of X-ray reflection nebulae. A&A 589, pp. A88. External Links: Document, 1511.05125, ADS entry Cited by: footnote 8.
- Discovery of Variable Iron Fluorescence from Reflection Nebulae in the Galactic Center. ApJL 656 (2), pp. L69–L72. External Links: Document, astro-ph/0611651, ADS entry Cited by: §1.
- Chandra/HETGS Observations of the Brightest Flare Seen from Sgr A*. ApJ 759 (2), pp. 95. External Links: Document, 1209.6354, ADS entry Cited by: §1.
- A powerful flare from Sgr A* confirms the synchrotron nature of the X-ray emission. MNRAS 468 (2), pp. 2447–2468. External Links: Document, 1703.03410, ADS entry Cited by: §1.
- Discovery of a Superluminal Fe K Echo at the Galactic Center: The Glorious Past of Sgr A* Preserved by Molecular Clouds. ApJ 714 (1), pp. 732–747. External Links: Document, 1003.2001, ADS entry Cited by: §1, §4, §6.
- Traces of Past Activity in the Galactic Centre. In Cosmic Rays in Star-Forming Environments, D. F. Torres and O. Reimer (Eds.), Astrophysics and Space Science Proceedings, Vol. 34, pp. 331. External Links: Document, 1210.3034, ADS entry Cited by: §1, §1, §4, §6, §6.
- Hard X-ray view of the past activity of Sgr A* in a natural Compton mirror. A&A 425, pp. L49–L52. External Links: Document, astro-ph/0408190, ADS entry Cited by: §1.
- Discrete sources as the origin of the Galactic X-ray ridge emission. Nature 458 (7242), pp. 1142–1144. External Links: Document, 0904.4649, ADS entry Cited by: §4.3, §4.
- X-Ray Echo from the Sagittarius C Complex and 500-year Activity History of Sagittarius A*. PASJ 65 (2), pp. 33. External Links: Document, 1211.4529, ADS entry Cited by: §1.
- The Galactic LMXB Population and the Galactic Centre Region. New A Rev. 88, pp. 101536. External Links: Document, 2006.05063, ADS entry Cited by: §6.
- SRG/ART-XC Galactic Bulge deep survey. II. Catalogue of point sources. MNRAS 529 (2), pp. 941–952. External Links: Document, ADS entry Cited by: §7.
- Polarized light from Sagittarius A* in the near-infrared Ks-band. A&A 576, pp. A20. External Links: Document, 1411.0006, ADS entry Cited by: §6.
- A polarized view of the hot and violent universe. ExA 51 (3), pp. 1109–1141. External Links: Document, ADS entry Cited by: §7.
- 25 years of XMM-Newton observations of the Sgr A complex: 3D distribution and internal structure of the clouds. A&A 695, pp. A52. External Links: Document, 2501.09737, ADS entry Cited by: §1, §1.
- Scattering of X-ray emission lines by neutral and molecular hydrogen in the solar atmosphere and in the vicinity of active galactic nuclei and compact X-ray sources. AstL 22 (5), pp. 648–663. External Links: Document, astro-ph/9608114, ADS entry Cited by: §4, footnote 8.
- The Center of the Galaxy in the Recent Past: A View from GRANAT. ApJ 407, pp. 606. External Links: Document, ADS entry Cited by: §1.
- Equivalent width, shape and proper motion of the iron fluorescent line emission from molecular clouds as an indicator of the illuminating source X-ray flux history. MNRAS 297 (4), pp. 1279–1291. External Links: Document, astro-ph/9805038, ADS entry Cited by: §7.
- X-Ray Imaging and Spectroscopy Mission. PASJ. External Links: Document, ADS entry Cited by: §7.
- An X-ray survey of the central molecular zone: Variability of the Fe K emission line. A&A 612, pp. A102. External Links: Document, 1712.04232, ADS entry Cited by: §1, §4, §6.
- The K-Alpha Lines in the Background X-Ray Spectrum and the Interstellar Gas in Galaxies. SvAL 6, pp. 353–356. External Links: ADS entry Cited by: §1.
- Chandra Cluster Cosmology Project. II. Samples and X-Ray Data Reduction. ApJ 692 (2), pp. 1033–1059. External Links: Document, 0805.2207, ADS entry Cited by: §3.2.
- The Imaging X-Ray Polarimetry Explorer (IXPE): Pre-Launch. JATIS 8 (2), pp. 026002. External Links: Document, 2112.01269, ADS entry Cited by: §1.
- Chandra X-ray Observatory (CXO): overview. In X-Ray Optics, Instruments, and Missions III, J. E. Truemper and B. Aschenbach (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 4012, pp. 2–16. External Links: Document, astro-ph/0004127, ADS entry Cited by: §3.2.
- Hard X-Ray Morphological and Spectral Studies of the Galactic Center Molecular Cloud Sgr B2: Constraining Past Sgr A* Flaring Activity. ApJ 815 (2), pp. 132. External Links: Document, 1507.08740, ADS entry Cited by: §1.
- Sgr A* flares: tidal disruption of asteroids and planets?. MNRAS 421 (2), pp. 1315–1324. External Links: Document, 1110.6872, ADS entry Cited by: §6.
Appendix A Regions of interest
Here we present a table that summarizes the parameters of the regions of interest used for the signal and (astrophysical) background estimations throughout the paper. This refers to the light curves in Sect. 4 and the spectral and polarization analysis in Sect.5.
| Region | RA,∘ | Dec,∘ | ,∘ | ,∘ | |
| Arches | 266.460 | 28.822 | 0.123 | 0.018 | 30 |
| Background | 266.628 | 28.801 | 0.217 | 0.097 | 90 |
| G0.13-0.11 (nt) | 266.592 | 28.882 | 0.132 | 0.112 | 60 |
| Sgr A | 266.566 | 28.891 | 0.112 | 0.097 | 270 |
| Sgr A (re) | 266.550 | 28.891 | 0.105 | 0.085 | 60 |
| Sgr A (bg) | 266.579 | 28.849 | 0.154 | 0.085 | 60 |
Appendix B Systematic uncertainties